OpenStudy (anonymous):

Solve for x log base 8 (3x+6)=3

OpenStudy (anonymous):

write in equivalent exponential form \[\log_b(x)=y\iff b^y=x\]

OpenStudy (anonymous):

in your example \[\log_8(3x+6)=3\iff 3x+6=8^3\]

OpenStudy (anonymous):

Thanks!